# Wildfire Risk Prediction: A Review

### v3

@@ 7.4.1 Separated Spatial and Temporal Modeling Methods
In another study, Bhowmik et al. (2023) proposed a U-shaped LSTM network combining UNet and LSTM, using seven days of meteorological, environmental, vegetation, and geological data to predict wildfire risk. The U-shaped LSTM network first used a three-layer UNet encoder to extract semantic information, which was then flattened and input into an LSTM. The LSTM output was passed through a UNet decoder to produce end-to-end pixel-level predictions of wildfire heatmaps for the next 24 hours to two weeks. This model achieved over 97
@@ 7.4.2 Coupled Spatiotemporal Modeling Methods
Compared to the separated spatial and temporal modeling methods, event sequence prediction algorithms like ConvLSTM are also widely used. For instance, Burge et al. (2021) evaluated the feasibility of using ConvLSTM to simulate wildfire propagation dynamics on three simulated datasets with varying terrain, wind, humidity, and vegetation density distributions. The experiments demonstrated that ConvLSTM’s predictive capability surpassed that of CNN. In the study by Prapas et al. (2021), which compared traditional machine learning models (RF), time series prediction algorithms (LSTM), image classification methods (CNN), and spatiotemporal modeling methods (ConvLSTM) for next-day wildfire danger prediction across Greece, several key findings emerged. For the RF model, training data was constructed by averaging each driving factor for the 10 days preceding the prediction date at each spatial location, resulting in a dataset with dimensions $f\times 1$, where $f$ is the number of driving factors. The LSTM model used time series input, with data dimensions of $f\times t$, where $t$ represents the time series length (10 days). For the image classification task, the spatial data was constructed as spatial patches centered on the target point, with dimensions $f\times h\times w$, where $h$ and $w$ are the height and width of the image. Lastly, for the spatiotemporal prediction task using ConvLSTM, the dataset dimensions were $f\times h\times w\times t$. The results showed that LSTM and ConvLSTM performed better than RF and CNN. Evaluating the binary classification results, LSTM outperformed ConvLSTM overall, but when using the area under the ROC curve as the evaluation metric, ConvLSTM slightly surpassed LSTM, with AUC scores of 0.926 and 0.920, respectively.
@@ 7.4.3 Spatially Enhanced Coupled Spatiotemporal Modeling Methods
Furthermore, Ji et al. (2024), following a similar branch structure to Eddin et al. (2023) and Jiang et al. (2023), proposed a static location-aware model based on ConvLSTM. This model aimed to integrate global static features (sine and cosine values of longitude and latitude, along with climate variables) with dynamic variables (e.g., meteorological, human, and vegetation cover factors) to predict global wildfire probability on sub-seasonal to seasonal scales. In this model, global static features were first convolved and max-pooled, and the resulting feature maps were multiplied with climate variables for data fusion. SKNet was then employed to extract higher-order features, which were upsampled and merged with the static features generated by the encoder. Finally, upsampling and skip connections were used to restore the image size and supplement the spatial details of global static location variables. The dynamic and static variables were concatenated and input into ConvLSTM. The hidden state output from the last time step of ConvLSTM was combined with the static and global features, and a $1\times 1$ convolution was applied to produce the final wildfire probability prediction. Comparing accuracy, recall, F1 score, precision, and Kappa coefficient with different variants of vision Transformer and ConvLSTM demonstrated the effectiveness of incorporating global location information and climate variables for enhancing predictive capabilities.
Similarly, He et al. (2024) developed a next-day wildfire risk prediction model for eastern China by combining CNN, ConvLSTM, and vision Transformer models, also employing a branch structure. In this model, eight CNN modules were used to model eight meteorological variables, while another CNN module modeled terrain, vegetation, and anthropogenic factors from the last day. The sequence feature maps generated by the temporal meteorological variables were input into LSTM, and its output was merged with the other variables’ CNN outputs. Channel attention, spatial attention, and CNN were used to extract higher-order features. To emphasize the high-resolution structural features of vegetation and scattered small settlements, Landsat data were used to generate 30-meter monthly high-resolution NDVI products, and vision Transformer was applied to extract these structural features. The flattened outputs from the vision Transformer and CNN were concatenated and classified using an artificial neural network (ANN) binary classifier. Ablation studies demonstrated the effectiveness of incorporating attention mechanisms, high-resolution spatial structure features, and vision Transformer. Sensitivity analysis also confirmed that meteorological factors were the most important among wildfire drivers.
@@
7.5 Promising New Methods
@@
One potential solution is Mamba (Gu and Dao, 2024), a parameter-efficient model designed to handle long sequence modeling (Xu et al., 2024) such as videos or remote sensing imagery over long temporal scales. Mamba’s architecture reduces the number of parameters through optimized dimensionality reduction and hierarchical attention mechanisms, allowing it to focus on key spatial and temporal features. This makes Mamba particularly promising for constructing wildfire risk prediction models that require high-resolution spatial and temporal data over extended time periods. The architecture of Mamba is shown below.
Moreover, pre-trained models and foundation models offer new opportunities for building more generalizable and robust wildfire risk assessment models. In various fields, domain-specific pre-trained or foundation models have been developed. Examples include the Segment Anything series (Kirillov et al., 2023; Ravi et al., 2024) for image segmentation, and the DINO series (Caron et al., 2021; Oquab et al., 2024) for general vision tasks. In the remote sensing field, foundation models have been proposed for image generation, segmentation, classification, and retrieval tasks (Li et al., 2024a; Zhang et al., 2024b; Hong et al., 2024; Li et al., 2024b; Kuckreja et al., 2024). These models, compared to traditional deep learning-based segmentation, classification, and retrieval models, can handle multi-modal data and establish relationships between semantic reasoning and visual concept modeling (Li et al., 2024a).
Wildfire risk prediction studies frequently highlight the importance of weather and climate factors, and in this context, several foundational models, such as FourCastNet (Kurth et al., 2023), Pangu-Weather (Bi et al., 2023), and GraphCast (Lam et al., 2023), have been developed specifically for weather and climate forecasting.
These foundation models, initially designed for image processing or remote sensing applications, can be easily adapted for downstream tasks like wildfire risk prediction and detection. A key research question is how to integrate these weather and climate forecasting products to enhance short-term wildfire risk predictions and improve the feasibility of large-scale, long-term wildfire risk forecasting.
Moreover, multi-source remote sensing satellites, including the Landsat series, Sentinel series, MODIS Terra and Aqua, and various meteorological satellites, provide long time series, large spatial coverage, and multi-band Earth observation products. These products are foundational for building wildfire detection algorithms, historical wildfire datasets, and quantitative estimation methods for wildfire driving factors. They also provide a rich dataset for training deep learning models or wildfire risk prediction foundation models. Thus, leveraging these multi-source, multi-modal, long-term Earth observation data and foundation model methodologies could lead to the development of foundational models for wildfire risk prediction. Ultimately, this approach aims to achieve high temporal-spatial resolution, large-scale, and long-term wildfire risk forecasting.
@@ 8 Conclusion
Another crucial aspect of wildfire prediction models is the choice of algorithms. The most commonly used models include traditional machine learning models, wildfire danger rating systems, and deep learning models. While traditional machine learning models have been extensively reviewed in the literature, deep learning models, despite their growing attention, have lacked a systematic overview. In this paper, deep learning models have been categorized into three types: time series prediction, image segmentation and classification, and spatiotemporal prediction models. The review of these models revealed a growing trend towards the consideration of longer time series and the development of higher spatiotemporal resolution wildfire risk assessment models.
Looking ahead, we foresee the potential for more efficient architectures, such as Mamba, and more generalized pre-trained and foundational models to further enhance the performance and scalability of deep learning-based wildfire risk prediction systems.

### v4

@@ 7.4.1 Separated Spatial and Temporal Modeling Methods
In another study, Bhowmik et al. (2023) proposed a U-shaped LSTM network combining UNet and LSTM, using seven days of meteorological, environmental, vegetation, and geological data to predict wildfire risk. The U-shaped LSTM network first used a three-layer UNet encoder to extract semantic information, which was then flattened and input into an LSTM. The LSTM output was passed through a UNet decoder to produce end-to-end pixel-level predictions of wildfire heatmaps for the next 24 hours to two weeks. This model achieved over 97% accuracy in predicting large-scale wildfires in California, outperforming CNN models.
@@ 7.4.2 Coupled Spatiotemporal Modeling Methods
Compared to the separated spatial and temporal modeling methods, time-series prediction algorithms like ConvLSTM are also widely used. For instance, Burge et al. (2021) evaluated the feasibility of using ConvLSTM to simulate wildfire propagation dynamics on three simulated datasets with varying terrain, wind, humidity, and vegetation density distributions. The experiments demonstrated that ConvLSTM’s predictive capability surpassed that of CNN. In the study by Prapas et al. (2021), which compared traditional machine learning models (RF), time series prediction algorithms (LSTM), image classification methods (CNN), and spatiotemporal modeling methods (ConvLSTM) for next-day wildfire danger prediction across Greece, several key findings emerged. For the RF model, training data was constructed by averaging each driving factor for the 10 days preceding the prediction date at each spatial location, resulting in a dataset with dimensions $f\times 1$, where $f$ is the number of driving factors. The LSTM model used time series input, with data dimensions of $f\times t$, where $t$ represents the time series length (10 days). For the image classification task, the spatial data was constructed as spatial patches centered on the target point, with dimensions $f\times h\times w$, where $h$ and $w$ are the height and width of the image. Lastly, for the spatiotemporal prediction task using ConvLSTM, the dataset dimensions were $f\times h\times w\times t$. The results showed that LSTM and ConvLSTM performed better than RF and CNN. Evaluating the binary classification results, LSTM outperformed ConvLSTM overall, but when using the area under the ROC curve as the evaluation metric, ConvLSTM slightly surpassed LSTM, with AUC scores of 0.926 and 0.920, respectively.
@@ 7.4.3 Spatially Enhanced Coupled Spatiotemporal Modeling Methods
Furthermore, Ji et al. (2024b), following a similar branch structure to Eddin et al. (2023) and Jiang et al. (2023), proposed a static location-aware model based on ConvLSTM. This model aimed to integrate global static features (sine and cosine values of longitude and latitude, along with climate variables) with dynamic variables (e.g., meteorological, human, and vegetation cover factors) to predict global wildfire probability on sub-seasonal to seasonal scales. In this model, global static features were first convolved and max-pooled, and the resulting feature maps were multiplied with climate variables for data fusion. SKNet was then employed to extract higher-order features, which were upsampled and merged with the static features generated by the encoder. Finally, upsampling and skip connections were used to restore the image size and supplement the spatial details of global static location variables. The dynamic and static variables were concatenated and input into ConvLSTM. The hidden state output from the last time step of ConvLSTM was combined with the static and global features, and a $1\times 1$ convolution was applied to produce the final wildfire probability prediction. Comparing accuracy, recall, F1 score, precision, and Kappa coefficient with different variants of vision Transformer and ConvLSTM demonstrated the effectiveness of incorporating global location information and climate variables for enhancing predictive capabilities.
Similarly, He et al. (2024) developed a next-day wildfire risk prediction model for eastern China by combining CNN, ConvLSTM, and vision Transformer models, also employing a branch structure. In this model, eight CNN modules were used to model eight meteorological variables, while another CNN module modeled terrain, vegetation, and anthropogenic factors from the last day of each time step. The sequence feature maps generated by the temporal meteorological variables were input into LSTM, and its output was merged with the other variables’ CNN outputs. Channel attention, spatial attention, and CNN were used to extract higher-order features. To emphasize the high-resolution structural features of vegetation and scattered small settlements, Landsat data were used to generate 30-meter monthly high-resolution NDVI products, and vision Transformer was applied to extract these structural features. The flattened outputs from the vision Transformer and CNN were concatenated and classified using an artificial neural network (ANN) binary classifier. Ablation studies demonstrated the effectiveness of incorporating attention mechanisms, high-resolution spatial structure features, and vision Transformer. Sensitivity analysis also confirmed that meteorological factors were the most important among wildfire drivers.
@@
7.5 Promising Deep Learning New Methods
Currently, the number and variety of features used in most wildfire risk prediction models have reached a plateau. However, due to the nonlinear and complex spatiotemporal relationships between various factors influencing wildfire ignition and spread, the trend in developing deep learning models for wildfire prediction focuses on enhancing their ability to model long-range spatiotemporal semantic dependencies. This approach aims to capture these subtle relationships more effectively. Specifically, there is a progression from time-series prediction to spatiotemporal sequence prediction and further to spatially enhanced spatiotemporal sequence prediction. Efforts are being made to improve the model’s spatiotemporal scale and resolution by incorporating higher-resolution geospatial data over extended time sequences.
With the rapid advancement of computing hardware, particularly GPUs, the scale of input data for deep learning models has continued to expand. However, current wildfire risk prediction models using RNNs, CNNs, Transformers, or GNNs are still limited by their efficiency and capacity to model large spatiotemporal scales, especially when compared to newer models like Mamba (Gu and Dao, 2024). Moreover, many of these models rely on multi-branch networks that use only images or nodes to capture long-range semantic dependencies, neglecting the potential efficiency gains offered by multimodal approaches, such as incorporating textual descriptions. This section will introduce recent advancements in deep learning for image segmentation tasks using multimodal techniques, followed by an overview of the parameter-efficient Mamba model.
7.5.1 Multimodal Models
Multimodal learning is a universal approach to building AI models, enabling the extraction and association of information from various sources, such as different sensors (Baltrušaitis et al., 2018). Unlike unimodal image-based learning, multimodal learning can acquire, interpret, and reason across multiple information sources, resembling human perception (Xu et al., 2023a). A key advantage of multimodal learning over unimodal approaches is its ability to more easily embed domain knowledge or abstract, knowledge-driven perspectives (Baltrušaitis et al., 2018). Currently, multimodal learning has been extensively studied in image segmentation and classification (Roy et al., 2023; Yao et al., 2023; Nasir et al., 2023; Chng et al., 2024; Wang and Ke, 2024; Xu et al., 2023b), data fusion (Khan et al., 2023; Kalamkar et al., 2023), generation, and visual-language pretraining models (Liu et al., 2023; Li et al., 2023a; Oquab et al., 2024). Given that this research focuses on exploring the potential applications of segmentation and classification methods in wildfire risk prediction, we will delve into Referring Image Segmentation (RIS) in subsequent sections.
RIS refers to models that understand natural language descriptions and correspond them to specific objects in images at the pixel level, outputting an object mask (Wu et al., 2024). RIS requires processing both text and visual modalities, with the latter typically being high-dimensional and exhibiting strong spatial characteristics encapsulating attributes such as color, texture, shape, and size. In contrast, natural language is generally low-dimensional and continuous, abstractly describing these visual features (Ji et al., 2024a). The key challenge of RIS lies in designing appropriate multimodal fusion methods to establish a relationship between language and visual modalities, thereby bridging the modality gap between the two (Chng et al., 2024). Generally, fusion methods can be categorized into early fusion and late fusion, commonly known as decoder-fusion and encoder-fusion, respectively (Wu et al., 2024).
Decoder fusion extracts visual and language features separately using different encoders, such as CNNs for visual data and LSTMs for language data, and designs specific fusion modules to combine these features at the decoder stage. For example, (sweere2022deep) developed Bottom-Up Shift and Bidirectional Attentive Refinement modules to progressively fuse the output features of the language and image modules. Similarly, Jain and Gandhi (2022) utilized CNNs to extract multi-stage image features and LSTMs for text feature extraction, integrating these features through a hierarchical cross-modal aggregation module. In another approach, Wang et al. (2022b) designed a knowledge transfer method to adapt image-level knowledge from Contrastive Language-Image Pretraining (CLIP) into RIS for pixel-level segmentation tasks, implementing a Neck module to fuse outputs from both the text and image encoders. Similarly, Yan et al. (2023) introduced a Fusion Neck module to combine image and text encoding information, alongside a multi-query generator module to aid segmentation.
Encoder fusion, in contrast, integrates multimodal features at the encoder stage using auxiliary losses or alignment modules (Chng et al., 2024). This method is considered more effective (Feng et al., 2021; Yang et al., 2022b) and has gained traction in recent studies (Wu et al., 2024). For example, Xu et al. (2023b) introduced a Bridger module, based on self-attention and cross-attention mechanisms, to facilitate cross-modal information exchange between each level of the text encoder and image encoder. Furthermore, the study designed a task decoder with hierarchical and global alignment to further fuse multimodal information and generate the mask prediction. Similarly, Yuan et al. (2024) employed BERT (Devlin et al., 2019) for text feature extraction and Swin Transformer (Liu et al., 2021b) for image feature extraction, combining them with a self-attention-based language-guided cross-scale enhancer and a pixel-word attention module for feature fusion. Moreover, Shah et al. (2024) incorporated a Gaussian Enhancement Module between textual and visual encoders to extract both local and global visual-language relationships, enhancing overall representation. In a similar vein, Chng et al. (2024) designed a Cross-modal Alignment Module between multiple layers of text and image encoders, addressing the granularity mismatch between the two modalities through bidirectional interactions. Additionally, they introduced a mask encoder branch and a Transformer-based masked token decoder to predict the mask.
In summary, multimodal tasks offer greater flexibility and intuitive interaction by incorporating natural language to guide the model, compared to unimodal tasks. Multimodal information provides additional contextual data that unimodal models often lack, allowing the language encoder to construct a relation graph for numerical geospatial relations (Yuan et al., 2024). Therefore, multimodal models, especially RIS, hold significant potential in wildfire risk prediction by capturing subtle spatiotemporal semantic dependencies that are difficult to model using only images or geospatial data.
7.5.2 Mamba
@@
One potential solution is Mamba: linear time sequence modeling with selective state spaces (Gu and Dao, 2024), a parameter-efficient model designed to handle long sequence modeling (Xu et al., 2024) such as videos or remote sensing imagery over long temporal scales. Mamba’s proposal is primarily aimed at addressing the computational complexity found in prevalent transformer-based networks. Specifically, the self-attention mechanism in transformers involves a computational complexity of $O(n^{2})$, where $n$ represents the sequence length. Mamba reduces this complexity from $O(n^{2})$ to $O(n)$. Compared to the Transformer network, which uses the self-attention mechanism, the recurrent neural network (RNN)—a previously dominant approach to sequence modeling—suffers from short memory when handling long sequences. Moreover, while RNNs are efficient in the inference stage, they are less efficient during the training phase. Mamba is built on top of the State Space Model (SSM), which, similar to RNNs, is used for sequence modeling. The state-space model can be represented as:
State Equation: $x_{t}=Ax_{t-1}+Bu_{t}+w_{t}$ (23)
Observation Equation: $y_{t}=Cx_{t}+Du_{t}+v_{t}$ (24)
Where $x_{t}$ is the state vector at time $t$, $u_{t}$ is the control/input vector, $y_{t}$ is the observation/output vector, $A$ is the state transition matrix, $B$ is the input matrix, $C$ is the observation matrix, $D$ is the direct transmission matrix (optional, sometimes set to 0), and $w_{t}$ and $v_{t}$ are process and observation noise, respectively, often assumed to be Gaussian noise and were often removed from most SSM equations.
By revising the vanilla State Space Model (SSM), as shown in Equations 23 and 24, the Structured State Space Model for Sequences (S4) was proposed by Gu et al. (2021). Specifically, since the data used in language and image processing are discrete, the SSM needed to be adapted into a discretized version. The zero-order hold technique was applied to discretize the SSM. Furthermore, the SSM equations were reformulated into a combination of recurrent and convolutional structures. This allows the S4 models to be trained in a convolutional neural network (CNN) manner while being tested in a recurrent neural network (RNN) manner, thus improving efficiency in both the training and inference stages.
To address the short memory issue typically found in RNNs, S4 employs the HiPPO matrix to replace matrix $A$ in Equation 23. This modification enables the model to store long sequence information in a compact and sparse matrix format. However, one limitation of the S4 models is that the matrices $A$, $B$, and $C$ are not data-dependent, which restricts their adaptability — this property is known as Linear Time Invariance (LTI).
To enhance flexibility, a gate-like operation was introduced to matrix $A$, allowing it to become data-dependent in a parameter-efficient manner. Similarly, matrices $B$, $C$, and the time step were adapted to be data-dependent using linear projection functions. Since all parameters are now data-dependent or dynamic, Gu and Dao (2023) proposed a parallel scan algorithm to efficiently operate the model on GPUs, enabling SSM models to scale for parallel computation.
While Mamba demonstrated effectiveness, especially with longer sequence lengths, it still did not fully capitalize on the power of GPUs. To address this, Gu and Dao (2024) introduced Mamba2, which applies a state-space duality approach. This new model leverages GPU capabilities more effectively, further improving the computational efficiency while maintaining high performance in long sequence modeling.
@@ 8 Conclusion
Another crucial aspect of wildfire prediction models is the choice of algorithms. The most commonly used models include traditional machine learning models, wildfire danger rating systems, and deep learning models. While traditional machine learning models have been extensively reviewed in the literature, deep learning models, despite their growing attention, have lacked a systematic overview. In this paper, deep learning models were categorized into three types: time series prediction, image segmentation and classification, and spatiotemporal prediction models. The review of these models revealed a growing trend towards the consideration of longer time series and the development of higher spatiotemporal resolution wildfire risk assessment models.
Looking ahead, we foresee the potential for more efficient architectures, such as Mamba, and multimodal models to further enhance the performance and scalability of deep learning-based wildfire risk prediction systems.